# Rebuilding ROME : Resolving Model Collapse during Sequential Model Editing

### v1

@@ Abstract
Recent work on model editing using Rank-One Model Editing (ROME), a popular model editing method, has shown that there are certain facts that the algorithm is unable to edit without breaking the model. Such edits have previously been called disabling edits Gupta et al. (2024). These disabling edits cause immediate model collapse and limits the use of ROME for sequential editing. In this paper, we make two main contributions. Firstly, we show that model collapse with ROME only happens when making edits using the CounterFact dataset Meng et al. (2022a) and does not happen when using the zsRE dataset Levy et al. (2017). Secondly, we find that disabling edits are an artifact of the original implementation of ROME. With this paper, we provide a more stable implementation ROME, which we call r-ROME and show that we no longer observe model collapse when making large scale sequential edits with ROME.
@@
Akshat Gupta, Gopala Anumanchipalli UC Berkeley akshat.gupta@berkeley.edu
@@ 1 Introduction
Large language models (LLMs) are expensive to train and the knowledge contained in these models gets obsolete with time. Model editing or knowledge editing (Yao et al., 2023) has recently come out as a popular method to update knowledge in large language models (LLMs). Knowledge editing is done in two main ways - either by adding information in-context (Mitchell et al., 2022; Zhong et al., 2023; Cohen et al., 2023), or using methods that modify the parameters of underlying model (De Cao et al., 2021; Mitchell et al., 2021; Meng et al., 2022a, b; Tan et al., 2023). In this paper, we focus on one popular parameter-modifying model editing methods called ROME (Rank-One Model Editing) Meng et al. (2022a).
@@
While a lot of model editing approaches perform well when making singular edits, editing multiple facts in a model still remains a challenge for parameter-modifying model editing methods. One way to make multiple edits to the same model is through sequential editing Yao et al. (2023) - where we make a series of single edits to a model by modifying the parameters of the model after every edit. Recent works have started studying the effects of sequential editing and show a gradual decay in model’s abilities as more edits are made to the model (Meng et al., 2022b; Yao et al., 2023; Gu et al., 2024; Gupta et al., 2024). Additionally, it was found that one popular model editing technique called ROME Meng et al. (2022a) was prone to a sudden collapse of the model by a single edit (Gupta et al., 2024; Yang et al., 2024; Hu et al., 2024). This effect was first observed in Gupta et al. (2024) during sequential editing. The collapse included complete loss of downstream performance, inability to recall previously editing facts and loss of the ability to even get edited. Such facts were named disabling edits by Gupta et al. (2024) and were later independently observed by Yang et al. (2024); Hu et al. (2024). Text generation examples for models post-collapse from a disabling edit are shown in Figure 1.
@@
Figure 1: A typical generation example after a disabling edit is compared to a normal model edit using ROME. The bold and underlined part in the text is the prompt given to the model, which is followed by the text generated by the model.
@@
Disabling edits are detrimental for knowledge editing at scale. While a gradual model degradation is expected as we make sequential edits to a model Gupta et al. (2024), disabling edits lead to a sudden model collapse irrespective of when the disabling fact is edited, making sequential editing impossible. An example of this can be seen in Figure 3(a), where instead of allowing gradual model degradation when doing sequential editing like in Figure 4(b), the presence of disabling edits lead to a sudden and immediate model collapse.
(a) CounterFact - ROME
(b) zsRE - ROME
Figure 2: This figure shows distribution of edits along |Delta| and Normalized Entropy metric for edits using ROME on CounterFact and zsRE dataset. These edits are for GPT2-XL model but we see a similar distribution for GPT-J.
In this paper, we aim to find the source of these disabling edits. We first introduce two metrics for identifying disabling edits - generation entropy and the norm of matrix update. We plot edits made by ROME along these two dimensions. We then perform large scale editing of GPT2-XL Radford et al. (2019) and GPT-J Wang and Komatsuzaki (2021) using ROME on two popular model editing datasets - CounterFact Meng et al. (2022a) and zsRE (Levy et al., 2017; Mitchell et al., 2021; Meng et al., 2022a). We find that disabling edits only exist when editing facts from the CounterFact dataset and not the zsRE dataset. After long inquiry into the reasons behind model collapse, we decide to rebuild the ROME code-base ourselves. To our surprise, we find that the problem of model collapse goes away with our implementation of ROME and we are able to perform large-scale sequential model editing using the algorithm. With this paper, we share our new ROME code-base and invite researchers to use it for model editing. Our implementation of ROME, which we call r-ROME, can be found here.
@@ 2 Background
The focus of our paper is a model editing algorithm called ROME (Rank-One Model Editing) Meng et al. (2022a). ROME is a parameter-modifying model editing method, which means that it inserts knowledge into LLMs by changing their parameters. ROME falls under the "locate-and-edit" category for model editing methods, where knowledge is updated by first identifying the layers in the model corresponding to stored knowledge and then modifying the weights for the specific layers.
Facts are usually added in a key-value format, where a key is the vector representation of a query-phrase and the value is the vector representation of the target object. For example, when adding a new fact - "The president of USA is John Cena", the query-phrase here is "The president of USA is" and the target object is "John Cena". Editing in ROME is done using a pair of vectors - $(k_{e},v_{e})$ that represent a new fact being added. $k_{e}$, also called the key-vector is a vector representation of the query-phrase, and $v_{e}$, or the value-vector is the vector representation of the target object. The weights of the specific layer being edited in ROME are updated from $W_{0}$ to $\hat{W}$ by inserting a new fact $(k_{e},v_{e})$ using the following equation:
@@
where $\Delta$ is the update to the current weight matrix being edited such that the new fact represented by $(k_{e},v_{e})$ is incorporated.
Dataset Implementation Efficacy Generalization Locality Fluency Score
ES $\uparrow$ EM $\uparrow$ PS $\uparrow$ PM $\uparrow$ NS $\uparrow$ NM $\uparrow$ GE $\uparrow$ S $\uparrow$
ROME Original $99.92$ $99.68$ $96.29$ $71.58$ $75.8$ $10.25$ $621.96$ $89.32$
Ours $99.78$ $99.64$ $97.4$ $71.16$ $74.86$ $10.0$ $621.08$ $89.16$
Table 1: A comparison between the original implementation Meng et al. (2022a) and our implementation of ROME on standard model editing metrics for GPT2-XL. The above table is created using a random sample of 5000 edits from the CounterFact dataset. We find that our implementation performs equally as well as the original implementation of ROME.
2.1 Evaluating Model Editing
Model editing is usually evaluated along three metrics - reliability, generalization and locality. Reliability measures if a fact was successfully added in a model, generalization measures if the edited fact is recalled through paraphrases of the prompt used to edit the fact, whereas locality measures if editing of one fact affects other facts stored inside a model. We follow standard model editing metrics proposed in Meng et al. (2022a). We refer the reader to Yao et al. (2023); Meng et al. (2022a) for a more comprehensive review of model editing metrics. Additionally, we also evaluated the model on downstream task performance as proposed by (Gu et al., 2024; Gupta et al., 2024), which becomes especially important when making sequential edits to the same model. We follow Gupta et al. (2024) and evaluate the edited model on four tasks from the GLUE Wang et al. (2018) benchmark - sentiment analysis (SST2) Socher et al. (2013), paraphrase detection (MRPC) Dolan and Brockett (2005), natural language inference (NLI) (Dagan et al., 2005; Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009) and linguistic acceptability classification Warstadt et al. (2019) for doing downstream evaluation.
@@ 3 Experiments
3.1 Metrics to Identify Disabling Edits
@@
Disabling edits Gupta et al. (2024) are defined as singular knowledge edits made to the model that renders it useless, leading to sudden loss of ability to do downstream tasks or any kind of meaningful generation. Gupta et al. (2024) also showed one way of identifying disabling edits was the unusually large norm of the update matrix. In other words, $|\Delta|$ in equation 1 was unusually higher when compared to the $|\Delta|$ for normal edits. We will use a modified version of $|\Delta|$ as the first metric to identify disabling edits, where we normalize $|\Delta|$ by the number of elements in the matrix.
@@
Additionally, Figure 1 shows a typical example of model collapse where the model constantly repeats a single word. This repeated word is most often the edited answer inserted into the model. The simplest metric to identify such a model collapse is to calculate the entropy over the probability distribution of vocabulary elements of text generated from the model. For this, a probability distribution is calculated over the vocabulary of a sample generation consisting of ten generations, and is normalized by the vocabulary size to remove the effect of the size of vocabulary. This constrains the entropy between 0 and 1. If the model collapses as shown in Figure 1, we expected the normalized entropy to be small and concentrated around a handful of words.
Note that this metric is a simplification of the generation entropy metric proposed by Meng et al. (2022a), where they take a weighted average of the entropy over of vocabulary of bigrams and trigrams. We chose to calculate entropy over just the unigram vocabulary words as it is cheaper and does the job, while normalization helps comparing entropy values for generations of different vocabulary size.
(a) Original Implementation
(b) Our Implementation
Figure 3: This figure shows the difference between the original updates and our implementation for GPT2-XL on the CounterFact dataset for 5k edits. As can be seen, our implementation significantly reduces potential disabling edits, with $|\Delta|$ being way lower.
3.2 Searching for Disabling Edits
The first set of experiments we do is to search for disabling edits. We do this by making large-scale singular model edits using ROME on GPT2-XL, on two popular model editing datasets - CounterFact Meng et al. (2022a) and zsRE (Levy et al., 2017; Mitchell et al., 2021). We measure the above mentioned metrics as shown in Figure 2.
We make the following two observations. Firstly, we observe that edits made using the zsRE dataset are very consistent along the two metrics, which means that the norm of the update are consistently small and generation entropy consistently high. The text generations made by all post-edit models when editing zsRE are consistently good and coherent, and do not have the repetitiveness found in disabling edits. We thus conclude that disabling edits are not present when editing facts using the zsRE dataset. Secondly, we see two clusters when editing facts from the CounterFact dataset. We find that certain edits have larger values of $|\Delta|$ for ROME, indicating the presence of disabling edits as seen in (Gupta et al., 2024; Yang et al., 2024; Hu et al., 2024). This shows that edits made using the CounterFact dataset can lead to model collapse. The above observation is quite surprising. It shows that there is a fundamental difference in the updates made to the model when editing facts using zsRE dataset and the CounterFact dataset. Model updates made using zsRE dataset have smaller $|\Delta|$ values and lead to healthy model behavior post editing.
@@ 3.2 Searching for Disabling Edits
Figure 5: Sequential editing with our re-implementation of ROME on GPT-J (6B).
The underlying reason for the difference in behavior of post-edit model is unclear, but is likely related to the characteristics of the two datasets. zsRE and CounterFact dataset differ in three major ways. Firstly, the CounterFact dataset contains counterfactual facts. This means that this dataset edits the model in such a way that lower probability objects are being inserted into the model, whereas in zsRE, we are inserting factually correct facts. The second difference in the two datasets is that zsRE edits facts using a question-answering prompt, whereas CounterFact prompts the model in a text completion format. This means that the key-vector in zsRE for the fact mentioned in section 2 corresponds to a question like "Who is the president of USA?" whereas for CounterFact it corresponds to "The president of USA is". Thirdly, all facts in CounterFact are one-word facts, are largely also tokenized into a single token for GPT2-XL and GPT-J, whereas most facts in zsRE contain multiple words. The difference in edits made by the two datasets is possibly due to one of the underlying reasons, but it is hard to answer this question without eliminating the confounding factors. Such a study is beyond the scope of this paper, and hence we continue to focus on removing disabling edits during model editing with ROME.
3.3 Fixing Disabling Edits
After a long inquiry into the optimization objective of ROME and studying its original code-base and different properties of the edits made by the algorithm, we found no reason for the $|\Delta|$ of certain edits to be so large. As a last resort, we re-implemented ROME ourselves. To our absolute surprise, we found that our re-implementation of ROME does not lead to disabling edits. The first evidence of this can be seen in Figure 3, where the $|\Delta|$ of the updates are orders of magnitude smaller for our implementation when compared to the original implementation. The comparison of original Meng et al. (2022a) and our implementation of ROME on standard model editing metrics is shown in Table 1. This table shows that our implementation performs equally well as original implementation of ROME, while having much smaller $|\Delta|$ values. We make no change to the update equation of ROME and re-implement it as is.
3.4 Sequential Editing with ROME
A large part of the reason why disabling edits went unnoticed initially was because large scale edits were not performed on the same model and the model was not evaluated on downstream tasks. Recent works (Gupta et al., 2024; Yang et al., 2024; Hu et al., 2024) observe disabling edits as a result performing sequential edits. Thus a final litmus test of our re-implementation of ROME is to perform sequential editing. The results for sequential editing can be seen in Figures 4 and 5.
Figure 4 shows a typical case of sequential editing using the original ROME code-base for GPT-J, where the presence of a disabling edit leads to large $|\Delta|$ and leads to model collapse, as can be seen by an immediate loss of downstream performance in Figure 3(a). With our re-implementation of ROME (Figure 5), we see that $|\Delta|$ is orders of magnitude smaller which allows the model to maintain its general abilities and avoids model collapse. This enables large scale sequential model editing without loss of performance. Similar trends can be seen for GPT2-XL in Figures 6 and 7.

### v3

@@ Abstract
Recent work using Rank-One Model Editing (ROME), a popular model editing method, has shown that there are certain facts that the algorithm is unable to edit without breaking the model. Such edits have previously been called disabling edits Gupta et al. (2024a). These disabling edits cause immediate model collapse and limits the use of ROME for sequential editing. In this paper, we show that disabling edits are an artifact of irregularities in the implementation of ROME. With this paper, we provide a more stable implementation ROME, which we call r-ROME and show that model collapse is no longer observed when making large scale sequential edits with r-ROME, while further improving generalization and locality of model editing compared to the original implementation of ROME.
@@
$\text{Akshat Gupta}^{1}$, $\text{Sidharth Baskaran}^{2}$, $\text{Gopala Anumanchipalli}^{1}$ ${}^{1}\text{UC Berkeley}$, ${}^{2}\text{Automorphic Inc.}$. akshat.gupta@berkeley.edu, sid@automorphic.ai
@@ 1 Introduction
Large language models (LLMs) are expensive to train and the knowledge contained in these models gets obsolete with time. Model editing or knowledge editing (Yao et al., 2023) has recently come out as a popular method to update knowledge in large language models (LLMs). In this paper, we focus on one popular parameter-modifying model editing methods called ROME (Rank-One Model Editing) Meng et al. (2022a). ROME is not only one of the most popular model editing algorithms, but is also widely used in unlearning Patil et al. (2023) and model interpretability (Ghandeharioun et al., 2024; Geva et al., 2023) literature.
@@
While a lot of model editing approaches perform well when making singular edits, editing multiple facts in a model still remains a challenge for parameter-modifying model editing methods. One way to make multiple edits to the same model is through sequential editing Yao et al. (2023) - where we make a series of single edits to a model by modifying the parameters of the model after every edit. Recent works have started studying the effects of sequential editing and found that ROME Meng et al. (2022a) was prone to a sudden model collapse by a single edit (Gupta et al., 2024a; Yang et al., 2024; Hu et al., 2024). This effect was first observed in Gupta et al. (2024a) during sequential editing. The collapse included complete loss of downstream performance, inability to recall previously editing facts and loss of the ability to even get edited. Such facts were named disabling edits by Gupta et al. (2024a) and were later independently observed by Yang et al. (2024); Hu et al. (2024).
@@
Figure 1: A typical generation example after a disabling edit is compared to a normal model edit using ROME. The bold and underlined part in the text is input prompt.
@@
Disabling edits are detrimental for knowledge editing at scale. While a gradual model degradation is expected as we make sequential edits to a model Gupta et al. (2024a), disabling edits lead to a sudden model collapse irrespective of when the disabling fact is edited, making sequential editing impossible. An example of this can be seen in Figure 3(a), where instead of allowing gradual model degradation when doing sequential editing like in Figure 4, the presence of disabling edits lead to a sudden and immediate model collapse.
Dataset Implementation Efficacy Generalization Locality Score
ES $\uparrow$ EM $\uparrow$ PS $\uparrow$ PM $\uparrow$ NS $\uparrow$ NM $\uparrow$ S $\uparrow$
CF Original $99.92$ $99.68$ $96.29$ $71.58$ $75.8$ $10.25$ $89.32$
r-ROME $99.74$ $97.79$ $\mathbf{99.09}$ $70.86$ $\mathbf{80.62}$ $26.0$ $\mathbf{92.22}$
p-ROME $\mathbf{99.9}$ $99.36$ $97.04$ $63.01$ $80.0$ $5.74$ $91.42$
Table 1: The above represents model editing results for 5000 singular model edits made on GPT-J-6B from the CounterFact dataset (non-sequential).
In this paper, we aim to find the source of these disabling edits. We first introduce two metrics for identifying disabling edits - generation entropy and the norm of matrix update. We plot edits made by ROME along these two dimensions and show new ways of identifying disabling edits even when making singular edits. As we dig deeper into the optimization objectives and the codebase of ROME, we find that the disabling edits in ROME are a result of irregularities in the implementation of ROME, and not an artifact of the optimization objective. Specifically, disabling edits were caused due to the asymmetric usage of key-vectors in the update equation of ROME. With this paper, we share our new ROME code-base and invite researchers to use it for model editing. Our implementation of ROME, which we call r-ROME, can be found here [^1] https://github.com/scalable-model-editing/rebuilding-rome.
@@ 2 Background
Facts are usually added in ROME using key-value format, where a key is the vector representation of a query-phrase and the value is the vector representation of the target object. For example, when adding a new fact - "The president of USA is John Cena", the query-phrase here is "The president of USA is" and the target object is "John Cena". The key-vector is defined by Meng et al. (2022a) is the activation of the first linear layer in the MLP targeted by ROME:
$k^{\left(l^{*}\right)}(x)=\sigma\left(W_{fc}^{\left(l^{*}\right)}\gamma\left(a_{[x],i}^{\left(l^{*}\right)}+h_{[x],i}^{\left(l^{*}-1\right)}\right)+b_{fc}^{\left(l^{*}\right)}\right)$ (1)
Editing in ROME is done using a pair of vectors - $(k_{e},v_{e})$ that represent a new fact being added. $k_{e}$, also called the key-vector is a vector representation of the query-phrase, and $v_{e}$, or the value-vector is the vector representation of the target object. The weights of the specific layer being edited in ROME are updated from $W_{0}$ to $\hat{W}$ by inserting a new fact $(k_{e},v_{e})$ using the following equation:
@@
where $\Delta$ is the update to the current weight matrix being edited such that the new fact $(k_{e},v_{e})$ gets incorporated. Additionally, each key-vector in $k_{e}$ is not just the representation of a single prompt. To enhance generalization, Meng et al. (2022a, b) create the key-vector as an average representations over the query-phrase with random prefixes. This is done so that the represented key-vectors do not just represent one way to phrase the query-phrase and edits made using these representations can generalize over different paraphrases of the edited facts. The final key vector is found by averaging over $N$ random prefixes using the equation:
$k_{e}=\frac{1}{N}\sum^{N}_{i=1}k(x_{i}\oplus p)$ (3)
Here $k(x_{i}\oplus p)$ represents the key-vector corresponding to a prefix $x_{i}$ being concatenated with the original query-phrase $p$. Examples of prefixes added in ROME can be seen in Table 3. In this paper, we will refer to the averaged prefix representation of keys with $k_{e}$, whereas when the representation just consists of the original prompt, we will depict that with a superscript as $k^{o}_{e}$. The following equation explicitly differentiates between the two mathematically:
$k^{o}_{e}=k(p)$ (4)
Evaluating Model Editing.
Model editing is usually evaluated along three metrics - reliability, generalization and locality. Reliability represents if a fact was successfully added in a model and is measured using edit score (ES) and edit magnitude (EM) metrics. ES measures the portion of cases when an edited fact is more probable than the original fact post-editing, whereas EM measures the difference in the probability magnitudes of the edited and original facts. Generalization represents if the edited fact is recalled through paraphrases of the prompt used to edit the fact and is measured using paraphrase score (PS) and paraphrase magnitude defined similary as above for paraphases of the edited facts. Locality represents if editing of one fact affects other facts stored inside a model and is measured using neighborhood score (NS) and neighborhood magnitude (NM) on facts unrelated to the edited facts. The score metric is the harmonic mean of ES, PS and NS. We follow standard model editing metrics proposed in the original ROME paper Meng et al. (2022a). We refer the reader to Yao et al. (2023); Meng et al. (2022a) for a more comprehensive review of model editing metrics.
Additionally, we also evaluated the model on downstream task performance as proposed by (Gupta et al., 2024a), which becomes especially important when making sequential edits to the same model. We evaluate the edited model on four tasks from the GLUE Wang et al. (2018) benchmark - sentiment analysis (SST2), paraphrase detection (MRPC), natural language inference (NLI) and linguistic acceptability classification for doing downstream evaluation.
(a) ROME
(b) r-ROME
Figure 2: This figure shows the difference between the ROME and r-ROME updates on GPTJ (6B) for 5k individual edits. Our implementation shows much less potential disabling edits indicated by lower $|\Delta|$ values.
Dataset Implementation Efficacy Generalization Locality Score
ES $\uparrow$ EM $\uparrow$ PS $\uparrow$ PM $\uparrow$ NS $\uparrow$ NM $\uparrow$ S $\uparrow$
CF Original $62.43$ $11.23$ $59.12$ $7.49$ $52.05$ $-0.05$ $57.53$
r-ROME $97.92$ $72.14$ $96.23$ $54.97$ $59.52$ $0.16$ $80.20$
p-ROME $99.94$ $95.31$ $94.05$ $55.22$ $52.57$ $-1.54$ $75.64$
Table 2: We find that our implementations (r-ROME & and p-ROME) retains edit performance significantly more than the original implementation of ROME on standard model editing metrics for GPT-J-6B. We use the same 5k CounterFact examples from as Table 1 sequentially.
@@ 3 Experiments
3.1 Properties of Disabling Edits
@@
Disabling edits Gupta et al. (2024a) are defined as singular knowledge edits that lead to sudden loss of ability to do downstream tasks or any kind of meaningful generation. Gupta et al. (2024a) also showed one way of identifying disabling edits was the unusually large norm of the update matrix. In other words, $|\Delta|$ in equation 2 was unusually higher when compared to normal edits.222$|\Delta|=\|\Delta\|_{2}/N$ is the L2 norm of the update matrix normalized by the number of elements in the update matrix.
@@
Figure 1 shows a typical example of model collapse where the model constantly repeats a single word. The simplest metric to identify such a model collapse is to calculate the entropy over the probability distribution of vocabulary elements of text generated from the model. For this, a probability distribution is calculated over the vocabulary of a sample generation consisting of ten generations, and is normalized by the vocabulary size to remove the effect of the size of vocabulary. If the model collapses as shown in Figure 1, we expected the normalized entropy to be small and concentrated around a handful of words.
The first set of experiments we do is to search for disabling edits. We do this by making singular model edits using ROME on GPT-J and GPT2-XL using the CounterFact dataset to replicate the conditions where disabling edits occurred in prior work. We measure the above mentioned metrics as shown in Figure 2(a) for GPT-J. Similar patterns are observed for GPT2-XL and are shown in Figure 5 (appendix). When editing facts from the CounterFact dataset, we see two clusters forming. We find that certain edits have larger values of $|\Delta|$ for ROME, indicating the presence of disabling edits.
3.2 Fixing ROME
After finding signals of disabling edits while making singular edits, we perform sequential editing with ROME. Every iteration of sequential editing with ROME leads to model collapse similar to Figure 3(a). This collapse occurs at random points during the editing process at one of the facts that clustered away in Figure 2(a). After a long inquiry into the optimization objective of ROME, we found no reason for $|\Delta|$ of certain edits to be so large. We then turned to the implementation of ROME and found some interesting discrepancies. Although seemingly benign, these discrepancies eventually lead to disabling edits. The core reason behind disabling edits is that instead of implementing equation 2 as mentioned in the paper, the authors of ROME Meng et al. (2022a) implement the following equation for $\Delta$:
$\hskip 10.0pt\Delta_{imp}=(v_{e}-W_{0}\mathbf{k^{o}_{e}})\frac{k_{e}^{T}C_{0}^{-1}}{k_{e}^{T}C_{0}^{-1}\mathbf{k^{o}_{e}}}$ (5)
where $\Delta_{imp}$ represents the actual implementation of $\Delta$ in the code by Meng et al. (2022a), with the difference highlighted in bold. The difference in implementation and original derivation of ROME is the use of two different types of key vectors. Rather than using key-vectors that average over prefix prompts or $k_{e}$ (eq 3), the authors end up using $k^{o}_{e}$ (eq 4) is certain places in the update equation. We find that this asymmetry in usage of the key-vector causes disabling edits.
To fix this issue, we create homogeneity in the usage of the key-vectors. We first use $k_{e}$ everywhere in the update equation, an implementation we refer to as r-ROME. This is the correct implementation of ROME as originally intended by the authors of Meng et al. (2022a). We then use keys generated using only the original prompts or $k^{o}_{e}$ homogeneously in the update equation, referred to as p-ROME. This also tests the hypothesis that using a key-vector averaged over random prefixes can create more generalizable edits.
The first evidence of removal of disabling edits can be seen in Figure 2, where the $|\Delta|$ of the updates are orders of magnitude smaller for r-ROME when compared to the original implementation. The overall results for independent edits are shown in Table 1. We find that edits made using r-ROME create more generalized edits at the slight expense of efficacy, resulting in a higher total edit score than the original implementation. p-ROME leads to increased efficacy and worse generalization resulting in a slightly lower edit score. This shows that homogeneity in using key-vectors is crucial in making model edits.
@@ 3.2 Fixing ROME
Figure 4: Sequential editing with r-ROME on GPT-J.
3.3 Sequential Editing with r-ROME
The final litmus test of r-ROME is to study its performance during large scale sequential editing. Figure 3 shows a typical case of sequential editing using the original ROME code-base for GPT-J, where the presence of a disabling edit leads to large $|\Delta|$ and model collapse, as can be seen by an immediate loss of downstream performance in Figure 3(a). With r-ROME (Figure 4), we see that $|\Delta|$ is orders of magnitude smaller and increases smoothly, which allows the model to maintain its general abilities and avoids model collapse. This enables large scale sequential model editing without loss of performance. The final model editing metrics after 5000 sequential edits for GPT-J are shown in Figure 2, with r-ROME significantly outperforming the original implementation of ROME. Additional sequential editing results using p-ROME and GPT-XL can be found in section B.